OpenStudy (anonymous):

simplify: e^(1+ln(x))

OpenStudy (anonymous):

\[\huge e^{1+lnx}=e^1\times e^{lnx}=e+x \]

OpenStudy (anonymous):

that last part should be multiplication ^^^

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\huge ex \]

OpenStudy (anonymous):

yw...:)